OpenStudy (anonymous):

quadratic equation: 8y^2=2y+1

jhonyy9 (jhonyy9):

8y2-2y-1=0 y_1,_2=(2+/- sqrt(4+32))/16 = (2+/-6)/16= -4/16 and 8/16 = -1/4 and 1/2

OpenStudy (anonymous):

how did you get 2 for the first b i keep getting -4

jhonyy9 (jhonyy9):

no you write in the first step this equation in the standard form

OpenStudy (anonymous):

nvm thanks lol I just caught what I was doing wrong, I kept squaring the first b thanks

jhonyy9 (jhonyy9):

ok

jhonyy9 (jhonyy9):

yw